OpenStudy (anonymous):

6t^2+23t-4=0 how do you solve this problem?

OpenStudy (anonymous):

I'm using the ac method

OpenStudy (anonymous):

what do u want to solve for

OpenStudy (anonymous):

6t^2 + 23t - 4 =0 product of the two roots = 6x -4 = -24 sum of two roots = 23 am thinking of 24 and -1 6t^2 + 23t - 4 =0 breaking the middle term 6t^2 +24t - t - 4 = 0 factoring 6t (t+4) -1( t+4) = 0 (6t-1) (t+4) =0 therefore: 6t-1 = 0 and t = 1/6 or t+4 = 0 and t =-4 solution for t = (1/6, -4)

OpenStudy (anonymous):

I'm sorry I'm little confused how did you get your answer?

OpenStudy (anonymous):

Never I see how you got it.

OpenStudy (anonymous):

*mind

OpenStudy (anonymous):

good

OpenStudy (anonymous):

Thank you.